# A Game-Theoretic Framework for Distributed Load Balancing: Static and Dynamic Game Models

### v1

@@ Abstract
Motivated by applications in job scheduling, queuing networks, and load balancing in cyber-physical systems, we develop and analyze a game-theoretic framework to balance the load among servers in both static and dynamic settings. In these applications, jobs/tasks are often held by selfish entities that do not want to coordinate with each other, yet the goal is to balance the load among servers in a distributed manner. First, we provide a static game formulation in which each player holds a job with a certain processing requirement and wants to schedule it fractionally among a set of heterogeneous servers to minimize its average processing time. We show that this static game is a potential game and admits a pure Nash equilibrium (NE). In particular, the best-response dynamics converge to such an NE after $n$ iterations, where $n$ is the number of players. We then extend our results to a dynamic game setting, where jobs arrive and get processed in the system, and players observe the load (state) on the servers to decide how to schedule their jobs among the servers in order to minimize their averaged cumulative processing time. In this setting, we show that if the players update their strategies using dynamic best-response strategies, the system eventually becomes fully load-balanced and the players’ strategies converge to the pure NE of the static game. In particular, we show that the convergence time scales only polynomially with respect to the game parameters. Finally, we provide numerical results to evaluate the performance of our proposed algorithms under both static and dynamic settings.
{IEEEkeywords}
Games theory, Nash equilibrium, potential games, load balancing, job scheduling, distributed algorithms, best-response dynamics.
@@ 1 Introduction
Distributed and uncoordinated load balancing is challenging when a set of strategic agents compete for a set of shared resources. There is a large body of literature on load balancing, in which the proposed schemes can be either static or dynamic. In static load balancing schemes, agents use prior knowledge about the system, whereas dynamic load balancing methods make decisions based on the system’s real-time state.
Load balancing in strategic environments has been studied using game theory in the existing literature, where, in most applications, the system is often modeled as congestion games [1, 2]. In congestion games, each player’s utility is determined by the resources they select and the number of other players choosing the same resources. If all players choose the most beneficial resource, it becomes congested, and all players suffer from the congestion cost. As a result, agents may prefer to choose less profitable resources to avoid congestion. Moreover, uncoordinated competition over shared resources has also been analyzed through the lens of multi-armed bandit games [3, 4]. In such games, a set of independent decision-makers sequentially chooses among a group of arms with the goal of maximizing their expected gain, without having any prior knowledge about the properties of each arm.
One of the limitations of the aforementioned models is the assumption that subsequent rounds of these games are independent. In other words, the outcome of a given round has no effect on the outcome of future rounds [5]. However, there is a carryover effect between rounds in queuing systems, caused by the jobs remaining in the queues. In fact, the average wait time for each queue in each round depends not only on the number of tasks sent to it but also on the number of jobs remaining in the queue from previous rounds. Given these limitations and inspired by applications in job scheduling and queuing networks, we propose and analyze a game-theoretic framework to achieve load balancing between servers in both static and dynamic settings. In these scenarios, jobs are managed by self-interested players who are unwilling to coordinate; nonetheless, the objective is to distribute the load across servers in a balanced and decentralized manner.
We first introduce a static load balancing game model in which each player holds a job with a certain processing requirement and seeks to schedule it fractionally across a set of heterogeneous servers to minimize their average processing time. Next, to account for the carryover effect, we extend our model to a dynamic game setting, where jobs arrive and are processed in the system over time. In this setting, players observe the load (state) on the servers and decide how to schedule their jobs to minimize their cumulative average processing time. In both settings, we show that if players update their strategies using static or dynamic best-response at each state, the system eventually becomes load-balanced, and the players’ strategies converge to the pure Nash equilibrium (NE) of the underlying game. Notably, we show that the convergence time scales polynomially with respect to the game parameters in both settings, which is further justified using extensive numerical experiments.
@@ 1.1 Related Work
Studying distributed load balancing through a game-theoretic lens is a natural approach. Despite its potential, relatively few studies have applied game theory to the load balancing problem [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. In [38], the authors model the load balancing problem in heterogeneous systems as a non-cooperative, repeated finite game and identify a structure for the best-response policy in pure strategies. They later introduce a distributed greedy algorithm to implement the best-response policy. However, since their setting is stateless, the proposed algorithm does not necessarily perform well in dynamic games. Moreover, [39] models the load balancing problem as a cooperative game among players.
@@
The work [40] studies the distributed load balancing problem using multi-agent reinforcement learning methods. In this approach, each server simultaneously works on multiple jobs, and the efficiency with which the server handles a job depends on its capacity and the number of jobs it processes. The problem is modeled as a stochastic system where each player has access to local information. To evaluate server efficiency, each player maintains an efficiency estimator, a vector summarizing the server’s performance, which also serves as the player’s state. The players’ policies are then formulated as functions of these states. They compare the performance of their algorithm to other existing algorithms, such as non-adaptive algorithms—i.e., algorithms where players use servers in a predetermined manner—while revealing that naive communication between players does not necessarily improve the system’s overall efficiency.
@@
Krishnasamy et al. [41] study a variant of the multi-armed bandit problem with the carryover effect, which is suitable for queuing applications. They consider a bandit setting in which jobs queue for service, and the service rates of the different queues are initially unknown. The authors examine a queuing system with a single queue and multiple servers, where both job arrivals to the queue and the service rates of the servers follow a Bernoulli distribution. At each time slot, only one server can serve the queue, and the objective is to determine the optimal server scheduling policy for each time slot. They further evaluate the performance of different scheduling policies against the no-regret policy [^1] The policy that schedules the optimal server in every time slot..
@@
More recently, [5] studied the multi-agent version of the queuing system in [41]. Their setting differs from ours in the sense that, in our setting, each server has a separate queue and processes jobs with a fixed service rate. However, [5] describes a system where each player has a queue, and servers process jobs with a fixed time-independent probability. All unprocessed jobs are then sent back to their respective queues. Queues also receive bandit feedback on whether their job was cleared by their chosen server. One concern in such settings is whether the system remains stable, despite the selfish behavior of the learning players. This concern does not arise in settings with a centralized scheduler, as these systems remain stable if the sum of arrival rates is less than the sum of service rates. However, this condition does not guarantee stability in settings with strategic players. In [5], the authors show that if the capacity of the servers is high enough to allow a centralized coordinator to complete all jobs even with double the job arrival rate, and players use no-regret learning algorithms, then the system remains stable.A no-regret learning algorithm minimizes the expected difference in queue sizes relative to a strategy that knows the optimal server. They later show in [42] that stability can be guaranteed even with an extra capacity of $\frac{e}{e-1}\approx 1.58$, but only if players choose strategies that maximize their long-run success rate.
1.2 Contributions and Organization
@@
  • Static Game Model: We formulate a static game in which each player holds a job with specific processing requirements and aims to schedule it fractionally across a set of servers in order to minimize its average wait time. We show that this static game is a potential game, and hence admits a pure NE. Additionally, we show that the best-response dynamics converge to this pure NE after a number of iterations, where the number of iterations is equal to the number of players.
  • Dynamic Game Extension: To account for the carryover effect in queuing systems, we extend the static model to a dynamic game setting. In this extension, jobs arrive and are processed over time, and players update their scheduling strategies based on the current state of the servers. We show that if players use dynamic best-response strategies, the system eventually reaches a load-balanced state, and their strategies converge to the pure NE of the static game. Furthermore, we show that convergence occurs in a polynomial number of iterations.
@@
The remainder of the paper is organized as follows: In Section 2.1, we introduce the static game model, followed by its extension to a dynamic game setting in Section 2.2. We present our main results for the static game model in Section 3. In Section 4, we propose our algorithm for the dynamic game model and analyze its convergence properties. Section 5 presents numerical simulations to evaluate the performance of the proposed algorithms. Finally, Section 6 concludes the paper and discusses future research directions. Omitted proofs and other auxiliary results are deferred to the Appendix.
@@ 2 Problem Formulation
In this section, we first introduce the static load balancing game and subsequently extend it to a dynamic game setting.
@@ 2.1 Static Load Balancing Game
In the static load balancing game, there are a set $[m]=\{1,2,\ldots,m\}$ of $m$ servers with service rates $\mu_{j}>0$ $\forall j\in[m]$, and a set $[n]=\{1,2,\ldots,n\}$ of $n$ players. Each player $i\in[n]$ holds a job with processing requirement (length) $\lambda_{i}$. This job can be fractionally scheduled on different servers. [^2] Since there is a one-to-one correspondence between jobs and players, we often use the terms players and jobs interchangeably. We also define $\lambda_{\max}=\max_{i}\lambda_{i}$ and $\lambda_{\min}=\min_{i}\lambda_{i}$ to be the maximum and minimum length of jobs among all the players, respectively.
@@
We further assume that each server $j$ is initially occupied with a load of $s^{0}_{j}\geq 0$ jobs. A graphical illustration of the model is provided in Fig. 1. In this example, we consider three servers with initial loads $s_{1}^{0},s_{2}^{0},s_{3}^{0}$ and service rates $\mu_{1}=2$, $\mu_{2}=1.5$, and $\mu_{3}=1.0$, respectively. Two players are present, with job lengths $\lambda_{1}=3$ and $\lambda_{2}=5$. As shown in the figure, Player 1 has distributed 2 units of its job on Server 1, 0.5 units on Server 2, and the remaining 0.5 units on Server 3. Meanwhile, Player 2 schedules 1 unit of its job on Server 1 and the remainder on Server 3.
@@
[…]
@@ 2.2 Dynamic Load Balancing Game
After that, the state of server $j\in[m]$ changes according to
@@
where $(\cdot)^{+}=\max\{0,\cdot\}$. Finally, the goal for each player $i$ is to follow a certain scheduling policy $\pi_{i}=(a^{t}_{i},t=1,2,\ldots)$ to minimize its overall averaged cumulative cost given by
@@
We note that Assumption 1 is the stability condition and is crucial in the dynamic setting. Otherwise, one could always consider a worst-case instance where, regardless of any scheduling policy, the length of the waiting jobs in the server queues will grow unboundedly.
3 Main Theoretical Results for the Static Load Balancing Game
In this section, we analyze the static load balancing game and show that it belongs to the class of potential games, for which the best response dynamics are known to converge to a pure Nash equilibrium [43]. We then characterize the best response dynamics in a closed-form as a solution to a convex program. This closed-form solution will be used in the subsequent section to extend our results to the dynamic game setting.
Theorem 1
The static load balancing game is a potential game and admits a pure NE. In particular, the best response dynamics, in which players iteratively and sequentially redistribute their jobs over the servers by best responding to others’ strategies, converge to a pure NE.
Proof 3.2.
We show that the static load balancing game is an exact potential game [43], that is, we show that for any two action profiles $(a_{i},a_{-i})$ and $(a^{\prime}_{i},a_{-i})$ that only differ in the action of player $i$, we have

### v2

@@ Abstract
Motivated by applications in job scheduling, queuing networks, and load balancing in cyber-physical systems, we develop and analyze a game-theoretic framework to balance the load among servers in static and dynamic settings. In these applications, jobs/tasks are held by selfish entities that do not want to coordinate with each other, yet the goal is to balance the load among servers in a distributed manner. First, we provide a static game formulation in which each player holds a job with a specific processing requirement and wants to schedule it fractionally among a set of heterogeneous servers to minimize its average processing time. We show that this static game is a potential game with a pure Nash equilibrium (NE). In particular, the best-response dynamics converge to such an NE after $n$ iterations, where $n$ is the number of players. Additionally, we bound the price of anarchy (PoA) of the static game in terms of game parameters. We then extend our results to a dynamic game setting, where jobs arrive and get processed, and players observe the load on the servers to decide how to schedule their jobs. In this setting, we show that if the players update their strategies using dynamic best-response, the system eventually becomes fully load-balanced and the players’ strategies converge to the pure NE of the static game. In particular, we show that the convergence time scales only polynomially with respect to the game parameters. Finally, we provide numerical results to evaluate the performance of our proposed algorithms.
keywords:
Game theory, Nash equilibrium, potential games, dynamic games, multi-agent systems, distributed load balancing.
thanks: This work was supported by the Air Force Office of Scientific Research under grant FA9550-23-1-0107.
thanks: Fatemeh Fardno is with the Department of Electrical and Computer Engineering and Coordinated Science Laboratory, University of Illinois Urbana-Champaign, Urbana, IL 61801.thanks: Rasoul Etesami is with the Department of Industrial and Systems Engineering, Department of Electrical and Computer Engineering, and Coordinated Science Laboratory, University of Illinois Urbana-Champaign, Urbana, IL 61801.
@@ 1 Introduction
Distributed and uncoordinated load balancing is challenging when a set of strategic agents compete for a set of shared resources. There is a large body of literature on load balancing, in which the proposed schemes can be either static or dynamic. In static load balancing schemes [1, 2, 3, 4], agents use prior knowledge about the system, whereas dynamic load balancing methods [5, 6, 7, 8, 9] make decisions based on the system’s real-time state.
Load balancing in strategic environments has been studied using game theory in the existing literature, where, in most applications, the system is often modeled as congestion games [10, 11]. In congestion games, each player’s utility is determined by the resources they select and the number of other players choosing the same resources. If all players choose the most beneficial resource, it becomes congested, and all players suffer from the congestion cost. As a result, agents may prefer to choose less profitable resources to avoid congestion.
One of the limitations of the aforementioned models is the assumption that rounds are independent, i.e., the outcome of one round does not affect future rounds [6]. However, in queuing systems, jobs left in queues create a carryover effect, making each round’s average wait time dependent not only on newly assigned jobs but also on residual jobs. Motivated by these limitations and applications in job scheduling and queuing networks, we propose and analyze a game-theoretic framework for load balancing in both static and dynamic settings. Here, jobs are managed by self-interested players who do not coordinate, yet the goal is to achieve decentralized and balanced server loads.
We first introduce a static load balancing game model in which each player holds a job with a certain processing requirement and seeks to schedule it fractionally across a set of heterogeneous servers to minimize their average processing time. Next, to account for the carryover effect, we extend our model to a dynamic game setting, where jobs arrive and are processed in the system over time. In this setting, players observe the load (state) on the servers and decide how to schedule their jobs to minimize their cumulative average processing time. We also assume that, at each time step, only one player receives a job. In both settings, we show that if players update their strategies using static or dynamic best-response at each state, the system eventually becomes load-balanced, and the players’ strategies converge to the pure Nash equilibrium (NE) of the underlying game. Notably, we show that the convergence time scales polynomially with respect to the game parameters in both settings, which is further justified using extensive numerical experiments.
@@ 1.1 Related Work
Studying distributed load balancing through a game-theoretic lens is a natural approach. Despite its potential, relatively few studies have applied game theory to the load balancing problem [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. In [2], the authors model the load balancing problem in heterogeneous systems as a non-cooperative, repeated finite game and identify a structure for the best-response policy in pure strategies. They later introduce a distributed greedy algorithm to implement the best-response policy. However, since their setting is stateless, the proposed algorithm does not necessarily perform well in dynamic games. Additionally, [3] models the load balancing problem as a cooperative game among players. Unlike distributed load balancing algorithms without queue memory [2, 3], which rely on stateless updates, our model explicitly captures the state dynamics and shows that stability and convergence can be guaranteed in polynomial time, even under strategic behavior of the players.
@@
The work [4] studies the distributed load balancing problem using multi-agent reinforcement learning methods. In this approach, each server simultaneously works on multiple jobs, and the efficiency with which the server handles a job depends on its capacity and the number of jobs it processes. The problem is modeled as a stochastic system where each player has access to local information. To evaluate server efficiency, each player maintains an efficiency estimator, a vector summarizing the server’s performance, which also serves as the player’s state. The players’ policies are then formulated as functions of these states. They compare the performance of their algorithm to other existing algorithms, such as non-adaptive algorithms, i.e., algorithms where players use servers in a predetermined manner, while revealing that naive communication between players does not necessarily improve the system’s overall efficiency.
@@
Krishnasamy et al. [9] study a variant of the multi-armed bandit problem with the carryover effect, which is suitable for queuing applications. They consider a bandit setting in which jobs queue for service, and the service rates of the different queues are initially unknown. The authors examine a queuing system with a single queue and multiple servers, where both job arrivals to the queue and the service rates of the servers follow a Bernoulli distribution. At each time slot, only one server can serve the queue, and the objective is to determine the optimal server scheduling policy for each time slot. They further evaluate the performance of different scheduling policies against the no-regret policy [^1] The policy that schedules the optimal server at each time..
@@
More recently, [6] studied the multi-agent version of the queuing system in [9]. Their setting differs from ours in the sense that, in our setting, each server has a separate queue and processes jobs with a fixed service rate. However, [6] describes a system where each player has a queue, and servers process jobs with a fixed time-independent probability. All unprocessed jobs are then sent back to their respective queues. Queues also receive bandit feedback on whether their job was cleared by their chosen server. One concern in such settings is whether the system remains stable, despite the selfish behavior of the learning players. This concern does not arise in settings with a centralized scheduler, as these systems remain stable if the sum of arrival rates is less than the sum of service rates. However, this condition does not guarantee stability in settings with strategic players. In [6], the authors show that if the capacity of the servers is high enough to allow a centralized coordinator to complete all jobs even with double the job arrival rate, and players use no-regret learning algorithms, [^2] A no-regret learning algorithm minimizes the expected difference in queue sizes relative to a strategy that knows the optimal server. then the system remains stable. They later show in [7] that stability can be guaranteed even with an extra capacity of $\frac{e}{e-1}$, where $e\approx 2.718$ is the natural number, but only if players choose strategies that maximize their long-run success rate. Subsequently, [5] extends the model in [6] by allowing servers to hold a single packet for later service if it cannot be served immediately, and by removing the need to prioritize older packets. Compared to no-regret learning in queuing systems [5, 6, 7], which emphasizes learning behavior under adversarial and stochastic arrival processes, our framework adopts a strategic, game-theoretic approach, where agents optimize best responses against evolving queue states. We establish polynomial-time convergence under a deterministic update rule, rather than relying on regret minimization.
Our dynamic game fundamentally differs from repeated congestion games with discounting [44, 45], which assume that each round is independent of the others. First, in repeated homogeneous congestion games, all players sharing a resource receive the same reward, which decreases with increased usage [46]. In contrast, our model allows heterogeneity, where players using the same resource may receive different rewards. Second, we incorporate a carryover effect: queue states evolve across rounds based on their current state and the number of players choosing them. Rather than optimizing discounted payoffs, we focus on long-run system behavior and establish convergence guarantees based on average cost over time.
Finally, our work is also related to [47], where the authors introduce a new class of Markov congestion games, which model non-stationarity in congestion games. They consider the carryover effect, and at each time step $t$ and state $s$, the players participate in a static congestion game. However, the authors show that even with a centralized algorithm, computing a NE remains computationally intractable. Our setting differs in two key aspects: we assume asynchronous policy updates for players and deterministic state transitions. Moreover, in our model, players using the same resource can receive different rewards. Under these assumptions, we show that our algorithm converges to a load-balanced equilibrium state in polynomial time.
1.2 Contributions
@@
  • Static Game Model: We formulate a static game in which each player holds one job with a specific processing requirement and aims to schedule it fractionally across a set of servers to minimize its average wait time. We show that this static game is a potential game and, therefore, admits a pure NE. Moreover, we show that best-response dynamics converge to this pure NE in a number of iterations equal to the number of players. Additionally, we bound the price of anarchy (PoA) of the static game in terms of game parameters, such as service rates, processing requirements, and initial loads.
  • Dynamic Game Extension: To account for the carryover effect in queuing systems, we extend the static model to a dynamic game setting, where at each time step, only one player receives a job. In this extension, jobs arrive and are processed over time, and players update their scheduling strategies based on the current state of the servers, which is fully observable to them. We show that if players use dynamic best-response strategies, the system eventually reaches a load-balanced state, and their strategies converge to the pure NE of the static game. Furthermore, we show that convergence occurs in a polynomial number of iterations.
@@
In summary, our work fills a gap in the literature by deriving a closed-form best response for both the static and dynamic load balancing games, proving single-pass convergence for strategic updates in the static game and bounding its price of anarchy, and establishing polynomial-time convergence for the dynamic game with queue carryover.
1.3 Organization and Notations
In Section 2, we introduce the static load-balancing game model, followed by its extension to a dynamic game setting. Section 3 presents our main results for the static game model. In Section 4, we propose an algorithm for the dynamic game model and analyze its convergence properties. Section 5 provides numerical simulations to evaluate the performance of the proposed algorithms. Finally, Section 6 concludes the paper and discusses directions for future research. Omitted proofs and other auxiliary results are provided in the Appendix. The main notations used in this paper are listed in Table 1.
Notation Definitions
$n$ number of players
$m$ number of servers
$\lambda_{i}$ job length of player $i$
$\lambda_{\text{max}}$ maximum job length across all players
$\lambda_{\text{min}}$ minimum job length across all players
$\mu_{j}$ service rate of server $j$
$\mu_{\text{max}}$ maximum service rate across all servers
$\mu_{\text{min}}$ minimum service rate across all servers
$a^{t}_{ij}$ fraction of job $i$ scheduled on server $j$ at time $t$
$a^{t}_{i}$ action vector of player $i$ at time $t$
$s_{j}^{0}$ initial load on server $j$
$s_{j}^{t}$ load on server $j$ at time $t$
Table 1: List of notations
@@ 2 Problem Formulation
In this section, we first introduce the static load balancing game and then extend it to a dynamic game setting.
@@ 2.1 Static Load Balancing Game
In the static load balancing game, there are a set $m]=\{1,2,\ldots,m\}$ of $m$ servers with service rates $\mu_{j}>0$ $\forall j\in[m]$, and a set $n]=\{1,2,\ldots,n\}$ of $n$ players. Each player $i\in[n]$ holds a job with processing requirement (length) $\lambda_{i}$. This job can be fractionally scheduled on different servers. [^3] Since there is a one-to-one correspondence between jobs and players, we often use the terms players and jobs interchangeably. We also define $\lambda_{\max}=\max_{i}\lambda_{i}$ and $\lambda_{\min}=\min_{i}\lambda_{i}$ to be the maximum and minimum length of jobs across all the players, respectively.
@@
We further assume that each server $j$ is initially occupied with a load of $s^{0}_{j}\geq 0$ jobs. A graphical illustration of the model is provided in Fig. 1. In this example, we consider three servers with initial loads $s_{1}^{0},s_{2}^{0},s_{3}^{0}$ and service rates $\mu_{1}=2$, $\mu_{2}=1.5$, and $\mu_{3}=1$, respectively. Two players are present, with job lengths $\lambda_{1}=3$ and $\lambda_{2}=5$. As shown in the figure, Player 1 has distributed 2 units of its job on Server 1, 0.5 units on Server 2, and the remaining 0.5 units on Server 3. Meanwhile, Player 2 schedules 1 unit of its job on Server 1 and the remainder on Server 3.
@@
$\int_{0}^{\alpha\lambda_{i}}\frac{x+s_{j}^{0}}{\mu_{j}}\ dx=\frac{\alpha^{2}\lambda_{i}^{2}}{2\mu_{j}}+\frac{\alpha\lambda_{i}s_{j}^{0}}{\mu_{j}}=\alpha\lambda_{i}\left(\frac{\alpha\lambda_{i}}{2\mu_{j}}+\frac{s_{j}^{0}}{\mu_{j}}\right).$
@@ 2.2 Dynamic Load Balancing Game
After that, the state of server $j$ changes according to
@@
where $(\cdot)^{+}=\max\{0,\cdot\}$.
Finally, the goal for each player $i$ is to follow a certain scheduling policy $\pi_{i}=(a^{t}_{i},t=1,2,\ldots)$ to minimize its overall averaged cumulative cost given by
@@
We note that Assumption 1 is the standard stability condition from queuing theory [48] and is crucial in the dynamic setting. Otherwise, one could always consider a worst-case instance where, regardless of any scheduling policy, the length of the waiting jobs in the server queues will grow unbounded.
We conclude this section with an example that illustrates the updates in the dynamic load-balancing game.
Example 2.1.
Consider a simple setting with $m=2$ servers and $n=2$ players. Suppose player 1’s job length is $\lambda_{1}=1$, and player 2’s job length is $\lambda_{2}=2$. The initial server loads are $s^{0}=(2,4)$, and the servers’ service rates are $\mu_{1}=1.5$ and $\mu_{2}=2.5$. Suppose that at time $t=0$, player 2 receives a job and decides to split it evenly between the two servers, i.e., $a_{21}=a_{22}=\frac{1}{2}$. An illustrative diagram of the model at time $t=0$ is shown in Fig. 2.
Figure 2: Illustration of the dynamic game at time $t=0$.
Using (4) we can write
$s_{1}^{1}=(2+2\left(\frac{1}{2}\right)-1.5)^{+}=1.5,$
$s_{2}^{1}=(4+2\left(\frac{1}{2}\right)-2.5)^{+}=2.5.$
Therefore, we have $s^{1}=(1.5,2.5)$, and player 2 incurs an instantaneous cost of
$D_{2}(a^{0},s^{0})$ $=\lambda_{2}a^{0}_{21}\left(\frac{\lambda_{2}a^{0}_{21}}{2\mu_{1}}+\frac{s_{1}^{0}}{\mu_{1}}\right)$
$+\lambda_{2}a^{0}_{22}\left(\frac{\lambda_{2}a^{0}_{22}}{2\mu_{2}}+\frac{s_{2}^{0}}{\mu_{2}}\right)$
$=\left(\frac{1}{3}+\frac{2}{1.5}\right)+\left(\frac{1}{5}+\frac{4}{2.5}\right)$
$=3.47.$
Now suppose that at time $t=1$, player 1 receives a job and sends all of it to server 1. Fig. 3 illustrates the dynamic game at time $t=1$.
Figure 3: Illustration of the dynamic game at time $t=1$.
3 Theoretical Results for the Static Game
In this section, we analyze the static load balancing game and show that it belongs to the class of exact potential games, for which a pure NE (defined below) is guaranteed to exist [49].
Definition 3.2.
An action profile $a=(a_{i},a_{-i})$ is called a pure NE for the static load balancing game if
$D_{i}(a_{i},a_{-i})\leq D_{i}(a^{\prime}_{i},a_{-i})\ \ \forall i,a^{\prime}_{i}\in A_{i}.$
In fact, in exact potential games, best response dynamics—where players iteratively and sequentially redistribute their jobs across servers by best responding to others’ strategies—are known to converge to a pure NE [49]. Thus, we next proceed to characterize the best response dynamics in a closed-form as a solution to a convex program. This closed-form solution will be used in the subsequent section to extend our results to the dynamic game setting. Our main results for the static load balancing game are presented in Lemma 5 and Theorems 2, 7, and 8.
Theorem 2.
The static load-balancing game is an exact potential game, that is, for any player $i$ and any two action profiles $(a_{i},a_{-i})$ and $(a^{\prime}_{i},a_{-i})$ that differ only in the action of player $i$, we have